# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v1

@@ 3.2 Discussion of the approach
The latent bottleneck structure and GRL have several important implications for inference, conditional generation, and uncertainty quantification. These implications are discussed here, and demonstrated through the case studies presented in Section 5. The main objective of the approach is to ensure that the known physics are properly utilized. Variability in the generative factors will be preferentially captured by the physics-grounded subset of the latent variables. The corresponding posterior distribution will be influenced by the domain and class contributions to the measured response, allowing for domain and class influences to be interpreted through their impact on the physics-grounded latent variables. A consequence of this behavior is that the posterior over physics-grounded latent variables will not necessarily be accurate, in the sense of a point estimate of a physical quantity obtained from the posterior being close to the underlying "true" value. This is also in part due to the data-driven component of the decoder, which yields a constant but not necessarily zero prediction when domain or class influences are not present in the measured response.
@@
The uncertainty associated with the predicted response stems from the approximate posterior distribution and the probabilistic decoder, and represents the combined influence of aleatoric and epistemic uncertainties. Without the GRL, and given sufficient data, the model would compensate for epistemic uncertainty in the form and parameters of the physics-based model in a data-driven manner. The estimated uncertainty in this case would only include the aleatoric component. In contrast, if no data-driven component is used in the decoder, the uncertainty would also include the epistemic uncertainty in the physics-based model. In our approach, part of the epistemic uncertainty due to domain and class influences will be accounted for in a data-driven manner. Therefore, the proposed approach is expected to yield uncertainty bounds somewhere in-between these two extremes. It is also important to note that the additivity assumption in Equation 2.3 is unlikely to hold for many physical systems. The model is expected to perform sub-optimally in such cases, resulting in increased uncertainty in the predictions.
@@
Finally, it is important to note that the proposed architecture is similar to the Domain Invariant Variational Autoencoder (DIVA) proposed by [Ilse2020], but differs in three key aspects. First, DIVA is targeted towards invariant representation learning and domain generalization in a purely data-driven setting. Additionally, DIVA utilizes an independence assumption between $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. In our proposed approach, the independence assumption is unnecessary due the constraints imposed by the GRL which is the third major difference. The lack of an independence assumption facilitates the use of a single probabilistic encoder, potentially reducing the amount of trainable parameters in the model and allowing for more complex and expressive encoder formulations. In our experiments we did not observe a decrease in performance when using a single encoder for all latent variables, compared to using separate encoders for each subset of the latent space, when combined with GRL regularization.
@@ 3.3 Quantitative assessment of the disentanglement
Quantitatively assessing disentanglement is a challenging problem and several metrics have been proposed [Higgins2016, Chen2018, Kim2018]. In practice, the degree of disentanglement achieved by a model is often evaluated based on subjective expectations stemming from domain knowledge, as noted in [Vowels2019]. Furthermore, recent work has demonstrated that the actual degree of disentanglement achieved by a particular model might have less influence on the value of certain metrics than the randomness in the initialization of the model parameters [Locatello2019], casting doubt on their performance in quantifying disentanglement. Our proposed approach aims to achieve "one-way" disentanglement and is conceptually more akin to techniques that temper the influence of misspecified model components [Carmona2020, Yu2022]. Intuitively, variations in generative factors that are described by the known physics should not affect the data-driven subsets of the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, while variations in generative factors not included in the known physics should still be preferentially captured by the physics-grounded subset of the latent variables. The degree to which this is achieved can be evaluated by comparing the amount of information captured by each subset $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ about a specified generative factor. When a subset of the latent variables is informative about a generative factor, it should be possible to train a regressor to predict the value of the generative factor from samples drawn from this subset of the latent variables. Based on this, we propose the following procedure to assess the amount of information about a given generative factor that is encoded in a subset of the latent variables for the trained model:
@@
This procedure yields $N_{f}$ sets of pair-wise $R^{2}$ values $\{R^{2}_{\bm{z}_{\mathrm{x}}\rightarrow s_{j}},R^{2}_{\bm{z}_{\mathrm{c}}\rightarrow s_{j}},R^{2}_{\bm{z}_{\mathrm{y}}\rightarrow s_{j}}\}_{j=1}^{N_{f}}$ with each of the $N_{f}$ sets corresponding to a single generative factor. A more informative subset of the latent variables should yield a more accurate regressor than an uninformative subset, and therefore also a higher $R^{2}$ value. It is emphasized that the metric described here is only intended to be a surrogate quantity for the amount of information encoded in each subset of the latent variables, and not a metric of disentanglement. Furthermore it requires access to the ground truth distribution and data generating process, and is therefore not generally applicable.
@@ 4 Previous work
Unsupervised disentanglement typically involves regularizing the VAE objective shown in Equation 2.1 by adjusting the relative importance of the KLD and reconstruction error terms. In the $\beta$-VAE architecture [Higgins2016], the KLD is scaled by a factor $\beta\geq 1$ that determines how much the approximate posterior is penalized for deviating from the prior distribution, limiting the capacity of the latent distribution and forcing independence between the latent variables at the expense of reconstruction quality. Alternative approaches involve weighting the importance of the total correlation term [Watanabe1960], a component of the KLD that quantifies and penalizes dependence between the dimensions of the aggregated posterior distribution (i.e. the posterior distribution marginalized over the entire dataset). To avoid the high computational cost associated with estimating the total correlation, recent works have proposed stochastic approximations based on mini-batches [Chen2018] and adversarial density-ratio estimation [Kim2018]. It has been shown that disentanglement can also be achieved by adjusting the MI between the input and the latent variables, leveraging the information bottleneck theory [Tishby2000]. In the variational setting, [Alemi2019] propose placing a constraint on the MI between the input and latent representation, and through Lagrange relaxation of the resulting constrained optimization problem obtain an objective that encourages the latent representation to act as a sufficient statistic of the input for reconstructing the output. The InfoVAE approach proposed by [Zhao2018] involves scaling certain terms in the ELBO, coupled with an additional term that promotes maximization of the MI between the inputs and latent variables. Several other approaches for unsupervised disentanglement have been proposed in the literature, such as enforcing independence between groups of latent variables [Lopez2018, Esmaeili2019], extending the standard architecture with adversarial components [Larsen2016], additional decoders [Ding2020] and sparsity inducing priors [Tonolini2020].
@@
Limited guidance is available on incorporating physics-based models in VAE, and particularly on the impact of combining NN components and physics-based models on the interpretability and generalizability of the predictions. In the context of modeling physical systems described by ordinary differential equations, the issue of balancing physics-based and data-driven components is highlighted by [Takeishi2021], and addressed using a regularized training objective based on the posterior predictive check [Gelman2013]. A similar setting is investigated in [Linial2021] and [Yildiz2019], with the latter introducing a regularized objective to ensure consistency of the latent space with the known physics. In [Walker2024] a VI approach is presented for discovering shared information in multi-modal data where the known physics are utilized to discover modality-dependent features that correlate with some part of the underlying physical process being studied. Recent work has explored the application of adversarial domain adaptation techniques to structural damage identification [Wang2022].
@@ 5 Synthetic case studies
Three synthetic case studies of different complexity, illustrated in Figure 1, are discussed in detail throughout this section. Each case addresses a different set of challenges. The Euler-Bernoulli beam case study demonstrates that the proposed approach preferentially utilizes the known physics, yielding an interpretable and parsimonious representation of the physical system. The damped oscillator case study highlights how the adversarial training prevents the model from learning arbitrary components of the response in a non-interpretable, data-driven manner. Finally, the bridge case study demonstrates the feasibility of using the model for damage detection in a more complex synthetic case, and compares the performance to that of existing data-driven approaches. For the purposes of reproducibility, the code needed to replicate the examples is made available on GitHub [^3] https://github.com/JanKoune/DPI-VAE.
@@
Three separate physics-based models are considered for every case study: A high-fidelity simulator, a full model, and a nominal model. The simulator is an accurate but generally computationally expensive model of the physical system, typically in the form of a Finite Element (FE) model, used to train the full and nominal models for each case study. The full model is a computationally efficient surrogate model of the ground truth data generating process: one or more structures with varying physical characteristics, subject to operational and environmental conditions, damage and degradation. To produce the training dataset for the full model, the simulator is evaluated on a set of generative factors $\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})\}_{i=1}^{N_{\mathrm{full}}}$, sampled uniformly and independently from prescribed ranges of values. The ranges are chosen to provide sufficient coverage over the support of the corresponding ground truth distribution $p_{\mathrm{gt}}(\bm{s}_{\mathrm{x}},\bm{s}_{\mathrm{c}},\bm{s}_{\mathrm{y}})$. The full model is then obtained by fitting a NN-based surrogate to the dataset composed of $N_{\mathrm{full}}$ input-output pairs $D_{\mathrm{full}}=\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}},\bm{x}^{(i)})\}_{i=1}^{N_{\mathrm{full}}}$ obtained from the simulator. Using a NN as the forward model for the data generating process enables the efficient visualization of the latent space and the reconstructions generated by the VAE for different inputs, simplifies the generation of test data to evaluate the performance of the VAE, and makes it possible to account for randomness in the hyperparameter initialization and data generation by averaging results over multiple runs with i.i.d. datasets. The nominal model corresponds to the available incomplete representation of the physics of the system under investigation. This could be setup independently, or starting from the simulator. In what follows, when an analytical expression describing the partially known physics is available, this is used as the nominal model. Alternatively, the nominal model is built by training a NN-based surrogate on a limited dataset $D_{\mathrm{nom}}=\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{x}^{(i)})\}_{i=1}^{N_{\mathrm{nom}}}$. The dataset $D_{\mathrm{nom}}$ is obtained by evaluating the simulator only on the physics-based subset of the generative factors $\{\bm{s}^{(i)}_{\mathrm{x}}\}_{i=1}^{N_{\mathrm{nom}}}$, while $\bm{s}_{\mathrm{c}}$ and $\bm{s}_{\mathrm{y}}$ are set to a constant reference value corresponding to the nominal condition of the structure.
@@
The VAE is trained, validated and tested on a dataset composed of $N_{\mathrm{total}}=N_{\mathrm{train}}+N_{\mathrm{val}}+N_{\mathrm{test}}$ triplets of observables $\mathcal{D}=\{(\bm{x}_{i},\bm{c_{i}},\bm{y_{i}})\}_{i=1}^{N_{\mathrm{total}}}$. This dataset is generated by first drawing samples of the generative factors from the ground truth distribution $(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})\sim{p_{\mathrm{gt}}(\bm{s}_{\mathrm{x}},\bm{s}_{\mathrm{c}},\bm{s}_{\mathrm{y}})}$ for $i=1,...,N_{\mathrm{total}}$, applying a set of deterministic transformations, and subsequently adding i.i.d. samples of zero-mean Gaussian white noise $\bm{\epsilon}_{\mathrm{x}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{x}}\bm{I})$, $\bm{\epsilon}_{\mathrm{c}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{c}}\bm{I})$ and $\bm{\epsilon}_{\mathrm{y}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{y}}\bm{I})$. Denoting the full model as $h_{\mathrm{x}}(\cdot)$, the response observables are obtained as $\bm{x}_{i}=h_{\mathrm{x}}(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})+\bm{\epsilon}_{\mathrm{x}}$. The domain and class observables are obtained as $\bm{c}_{i}=h_{\mathrm{c}}(\bm{s}^{(i)}_{\mathrm{c}})+\bm{\epsilon}_{\mathrm{c}}$ and $\bm{y}_{i}=h_{\mathrm{y}}(\bm{s}^{(i)}_{\mathrm{y}})+\bm{\epsilon}_{\mathrm{y}}$ respectively. For simplicity, $h_{\mathrm{c}}$ and $h_{\mathrm{y}}$ are taken as the identity function in all the case studies presented in this section. This procedure is illustrated in Figure 4. In all case studies we use $N_{\mathrm{train}}=1024$ and $N_{\mathrm{val}}=N_{\mathrm{test}}=512$.
@@
For all case studies there is no regularization considered other than the GRL, i.e. $\beta=\alpha_{\mathrm{x}}=\alpha_{\mathrm{c}}=\alpha_{\mathrm{y}}=1.0$. To enable the formulation of problems with bounded latent variables and to ensure a stable optimization procedure, the physics-based latent variables are internally transformed to an unbounded and normalized workspace. All physics-grounded latent variables are constrained to lie within ranges that ensure consistency with the underlying physics of each system. Additional details regarding the architecture, data, optimization, and plotting are provided in Appendix A. It is noted that the case studies are only meant as didactic examples, intended to elaborate the issues with combining physics-based and data-driven components in VAE, provide intuition about the interaction between these components, and demonstrate the behavior of the model. Therefore, emphasis is placed on clarity rather than realism.
5.1 Simply supported beam
The case study consists of a Euler-Bernoulli beam with fixed length $L=1.0$ m and a point load with magnitude $F=1.0$ N acting on an uncertain position $x_{\mathrm{F}}$ along the length of the beam. The material is linear elastic, with uncertain Young’s modulus $E$, area moment of inertia $I=2.0$ m4, and cross sectional area $A=1.0$ m2. The rotational stiffness of the right-hand side support is temperature dependent, with the dependence modeled as an increase in the rotational stiffness of the support at lower temperatures, which is gradually reduced to zero (corresponding to a pinned support) at higher temperatures. The relationship between the temperature and the support rotational stiffness is formulated as $\log k_{\mathrm{r}}=8-\frac{10}{1+e^{-T}}$. The beam is subject to variability in the vertical stiffness of the right-hand side support, e.g. due to damage or a deficiency of the support, simulated as a translational spring boundary condition with stiffness $k_{\mathrm{v}}$. This quantity can span several orders of magnitude, and therefore we parametrize the model using $\log k_{\mathrm{v}}$ instead. The beam is equipped with $d_{\mathrm{x}}=32$ displacement sensors, equally spaced along the length as shown in Figure 1(a). It is assumed that the sensors yield i.i.d. samples of the displacement field along the beam.

### v2

@@ 3.2 Discussion of the approach
The latent bottleneck structure and GRL have several important implications for inference, conditional generation, and uncertainty quantification. These are discussed here, and demonstrated through the case studies presented in Section 5. The main objective of the approach is to ensure that the known physics are properly utilized. Variability in the generative factors will be preferentially captured by the physics-grounded subset of the latent variables. The corresponding posterior distribution will be influenced by the domain and class contributions to the measured response, allowing for domain and class influences to be interpreted through their impact on the physics-grounded latent variables. Therefore, the posterior over physics-grounded latent variables might not necessarily be accurate in the sense of a point estimate of a physical quantity obtained from the posterior being close to the underlying “true value”. This is also in part due to the data-driven component of the decoder, which can yield a constant but not necessarily zero prediction when domain or class influences are not present in the measured response.
@@
The uncertainty associated with the predicted response stems from the approximate posterior distribution and the probabilistic decoder, and represents the combined influence of aleatoric and epistemic uncertainties. Without the GRL, and given sufficient data, the model would compensate for systematic discrepancies between the physics-based model and response measurements in a data-driven manner. In this case, the uncertainty in the reconstructed response would only represent aleatoric uncertainty. In contrast, if no data-driven component is used in the decoder, the uncertainty would also include epistemic uncertainty due to domain and class influences that are not included in the physics-based model. In our approach, part of this epistemic uncertainty is accounted for in a data-driven manner. Therefore, the proposed approach is expected to yield uncertainty bounds somewhere in-between these two extremes. We emphasize that the estimated uncertainty does not include the uncertainty over model parameters $\bm{\theta}$, $\bm{\phi}$ and $\bm{\psi}$. Finally, it is important to consider that the additivity assumption in Equation 2.3 is unlikely to hold for many physical systems. The model is expected to perform sub-optimally in such cases, resulting in inaccurate uncertainty estimates. None of the case studies presented in 5 satisfy the additivity assumption, demonstrating that the model can still be feasibly applied in such cases.
@@
Finally, it is important to note that the proposed architecture is similar to the Domain Invariant Variational Autoencoder (DIVA) proposed by [Ilse2020], but differs in three key aspects. First, DIVA is targeted towards invariant representation learning and domain generalization in a purely data-driven setting. Additionally, DIVA imposes an independence assumption between $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. In our proposed approach, the independence assumption is unnecessary due to the adversarial training, which is the third major difference. The lack of an independence assumption facilitates the use of a single probabilistic encoder, potentially reducing the amount of trainable parameters in the model and allowing for more complex and expressive encoder formulations. In our experiments we did not observe a decrease in performance when using a single encoder for all latent variables, compared to using separate encoders for each subset of the latent space, when combined with adversarial training.
@@ 3.3 Quantitative assessment of the disentanglement
Quantitatively assessing disentanglement is a challenging problem and several metrics have been proposed [Higgins2016, Chen2018, Kim2018]. This difficulty can be partially attributed to the lack of a consistent definition of disentanglement [Locatello2019]. In practice, the degree of disentanglement achieved by a model is often evaluated based on subjective expectations stemming from domain knowledge [Vowels2019]. Our proposed approach aims to achieve “one-way” disentanglement and is conceptually more akin to techniques that temper the influence of misspecified model components [Carmona2020, Yu2022]. Intuitively, variations in generative factors that are described by the known physics should not affect the data-driven subsets of the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, while variations in generative factors not included in the known physics should still be preferentially captured by the physics-grounded subset of the latent variables. The degree to which this is achieved can be evaluated by comparing the amount of information captured by each subset $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ about a specified generative factor. When a subset of the latent variables is informative about a generative factor, it should be possible to train a regressor to predict the value of the generative factor from samples drawn from this subset of the latent variables. Based on this, we propose the following procedure to assess the amount of information about a given generative factor that is encoded in a subset of the latent variables for the trained model:
@@
This procedure yields $N_{f}$ sets of pair-wise $R^{2}$ values $\{R^{2}_{\bm{z}_{\mathrm{x}}\rightarrow s_{j}},R^{2}_{\bm{z}_{\mathrm{c}}\rightarrow s_{j}},R^{2}_{\bm{z}_{\mathrm{y}}\rightarrow s_{j}}\}_{j=1}^{N_{f}}$ with each of the $N_{f}$ sets corresponding to a single generative factor. A more informative subset of the latent variables should yield a more accurate regressor than an uninformative subset, and therefore also a higher $R^{2}$ value. It is emphasized that the metric described here is only intended to be a surrogate quantity for the amount of information encoded in each subset of the latent variables, and not a metric of disentanglement. Furthermore, this metric requires access to the ground truth distribution and data generating process, and is therefore not generally applicable.
@@ 4 Previous work
Unsupervised disentanglement necessarily relies on inductive bias and implicit supervision [Locatello2019]. A common approach is to adjust the relative importance of the KLD and reconstruction error terms. In the $\beta$-VAE architecture [Higgins2016], the KLD is scaled by a factor $\beta\geq 1$ that determines how much the approximate posterior is penalized for deviating from the prior distribution, limiting the capacity of the latent distribution and forcing independence between the latent variables at the expense of reconstruction quality. Alternative approaches involve weighting the importance of the total correlation term [Watanabe1960], a component of the KLD that quantifies and penalizes dependence between the dimensions of the aggregated posterior distribution (i.e. the posterior distribution marginalized over the entire dataset). To avoid the high computational cost associated with estimating the total correlation, recent works have proposed stochastic approximations based on mini-batches [Chen2018] and adversarial density-ratio estimation [Kim2018]. It has been shown that disentanglement can also be achieved by adjusting the MI between the input and the latent variables, leveraging the information bottleneck theory [Tishby2000]. [Alemi2019] extend this approach to the variational setting, placing a constraint on the MI between the input and latent representation, and through Lagrange relaxation of the resulting constrained optimization problem obtain an objective that encourages the latent representation to act as a sufficient statistic of the input for reconstructing the output. The InfoVAE approach proposed by [Zhao2018] involves scaling specific terms in the ELBO, coupled with an additional term that promotes maximization of the MI between the inputs and latent variables. Several other approaches for unsupervised disentanglement have been proposed in the literature, such as enforcing independence between groups of latent variables [Lopez2018, Esmaeili2019], extending the standard architecture with adversarial components [Larsen2016], additional decoders [Ding2020] and sparsity inducing priors [Tonolini2020].
@@
Limited guidance is available on incorporating physics-based models in VAE, and particularly on the impact of combining NN components and physics-based models on the interpretability and generalizability of the predictions. In the context of modeling physical systems described by ordinary differential equations, the issue of balancing physics-based and data-driven components is highlighted by [Takeishi2021], and addressed using a regularized training objective based on the posterior predictive check [Gelman2013]. A similar setting is investigated in [Linial2021] and [Yildiz2019], with the latter introducing a regularized objective to ensure consistency of the latent space with the known physics. [Walker2024] present an approach for discovering shared information in multi-modal data where the known physics are utilized to discover modality-dependent features that correlate with some part of the underlying physical process being studied. Recent work has explored the application of adversarial domain adaptation techniques to structural damage identification [Wang2022].
@@ 5 Synthetic case studies
Three synthetic case studies of different complexity, illustrated in Figure 1, are discussed in detail throughout this section. Each case addresses a different set of challenges. The beam case study demonstrates that the proposed approach preferentially utilizes the known physics, yielding an interpretable and parsimonious representation of the physical system. The oscillator case study highlights how the adversarial training can prevent the model from learning arbitrary components of the response in a data-driven manner, and investigates the impact of the GRL hyperparameter. Finally, the bridge case study demonstrates the feasibility of using the model for damage detection in a more complex synthetic case, and compares the performance to that of existing data-driven approaches. For the purposes of reproducibility, the code needed to replicate the examples is made available on GitHub [^3] https://github.com/JanKoune/DPI-VAE.
@@
Three separate physics-based models are considered for every case study: A high-fidelity simulator, a full model, and a nominal model. The simulator is an accurate but generally computationally expensive model of the physical system, typically in the form of a Finite Element (FE) model, used to train the full and nominal models for each case study. The full model is a computationally efficient surrogate model of the ground truth data generating process: one or more structures with varying physical characteristics, subject to operational and environmental conditions, damage and degradation. To produce the training dataset for the full model, the simulator is evaluated on a set of generative factors $\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})\}_{i=1}^{N_{\mathrm{full}}}$, sampled uniformly and independently from prescribed ranges of values. The ranges are chosen to provide sufficient coverage over the support of the corresponding ground truth distribution $p_{\mathrm{gt}}(\bm{s}_{\mathrm{x}},\bm{s}_{\mathrm{c}},\bm{s}_{\mathrm{y}})$. The full model is then obtained by fitting a NN-based surrogate to the dataset composed of $N_{\mathrm{full}}$ input-output pairs $D_{\mathrm{full}}=\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}},\bm{x}^{(i)})\}_{i=1}^{N_{\mathrm{full}}}$ obtained from the simulator. Using a NN as the forward model for the data generating process enables the efficient visualization of the latent space and the reconstructions generated by the VAE for different inputs, simplifies the generation of test data to evaluate the performance of the VAE, and makes it possible to account for randomness in the hyperparameter initialization and data generation by averaging results over multiple runs with i.i.d. datasets. The nominal model corresponds to the available incomplete representation of the physics of the system under investigation. When an analytical expression describing the partially known physics is available, this is used as the nominal model. Alternatively, the nominal model is built by training a NN-based surrogate on a limited dataset $D_{\mathrm{nom}}=\{(\bm{s}^{(i)}_{\mathrm{x}},\bm{x}^{(i)})\}_{i=1}^{N_{\mathrm{nom}}}$, obtained by evaluating the simulator only on the physics-based subset of the generative factors $\{\bm{s}^{(i)}_{\mathrm{x}}\}_{i=1}^{N_{\mathrm{nom}}}$, while $\bm{s}_{\mathrm{c}}$ and $\bm{s}_{\mathrm{y}}$ are set to a constant reference value corresponding to the nominal condition of the structure.
@@
The VAE is trained and validated on a dataset composed of $N_{\mathrm{total}}=N_{\mathrm{train}}+N_{\mathrm{val}}$ triplets of observables $\mathcal{D}=\{(\bm{x}_{i},\bm{c_{i}},\bm{y_{i}})\}_{i=1}^{N_{\mathrm{total}}}$. This dataset is generated by first drawing samples of the generative factors from the ground truth distribution $(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})\sim{p_{\mathrm{gt}}(\bm{s}_{\mathrm{x}},\bm{s}_{\mathrm{c}},\bm{s}_{\mathrm{y}})}$ for $i=1,...,N_{\mathrm{total}}$, applying a set of deterministic transformations, and subsequently adding i.i.d. samples of zero-mean Gaussian white noise $\bm{\epsilon}_{\mathrm{x}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{x}}\bm{I})$, $\bm{\epsilon}_{\mathrm{c}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{c}}\bm{I})$ and $\bm{\epsilon}_{\mathrm{y}}\sim\mathcal{N}(\bm{0},\bm{\sigma}^{2}_{\mathrm{y}}\bm{I})$. Denoting the full model as $h_{\mathrm{x}}(\cdot)$, the response observables are obtained as $\bm{x}_{i}=h_{\mathrm{x}}(\bm{s}^{(i)}_{\mathrm{x}},\bm{s}^{(i)}_{\mathrm{c}},\bm{s}^{(i)}_{\mathrm{y}})+\bm{\epsilon}_{\mathrm{x}}$. The domain and class observables are obtained as $\bm{c}_{i}=h_{\mathrm{c}}(\bm{s}^{(i)}_{\mathrm{c}})+\bm{\epsilon}_{\mathrm{c}}$ and $\bm{y}_{i}=h_{\mathrm{y}}(\bm{s}^{(i)}_{\mathrm{y}})+\bm{\epsilon}_{\mathrm{y}}$ respectively. This procedure is illustrated in Figure 4.
@@
For all the case studies presented in this section, $h_{\mathrm{c}}$ and $h_{\mathrm{y}}$ are taken as the identity function for simplicity. Furthermore we use $N_{\mathrm{train}}=1024$ and $N_{\mathrm{val}}=512$, and consider no other regularization except for the GRL, i.e. $\beta=\alpha_{\mathrm{x}}=\alpha_{\mathrm{c}}=\alpha_{\mathrm{y}}=1.0$. To enable the formulation of problems with bounded latent variables and to ensure a stable optimization procedure, the physics-based latent variables are internally transformed to an unbounded and normalized workspace. All physics-grounded latent variables are constrained to lie within ranges that ensure consistency with the underlying physics of each system. Additional details regarding the architecture, data, optimization, and plotting are provided in Appendix A. It is noted that the case studies are only meant as didactic examples, intended to elaborate the issues with combining physics-based and data-driven components in VAE, provide intuition about the interaction between these components, and demonstrate the behavior of the model. Therefore, emphasis is placed on clarity rather than realism.
5.1 Beam case study
The case study consists of a beam with fixed length $L=1.0$ m and a point load with magnitude $F=1.0$ N acting on an uncertain position $x_{\mathrm{F}}$ along the length of the beam. The material is linear elastic, with uncertain Young’s modulus $E$, Poisson ratio $\nu=0.3$l area moment of inertia $I=2\cdot 10^{-6}$ m4 and $A=2.4\cdot 10^{-3}$ m2. The rotational stiffness of the right-hand side support is temperature dependent, with the dependence modeled as an increase in the rotational stiffness of the support at lower temperatures. The relationship between the temperature and the support rotational stiffness is formulated as $\log k_{\mathrm{r}}=8-\frac{10}{1+e^{-T/2}}$. The beam is subject to variability in the vertical stiffness of the right-hand side support, e.g. due to damage or a deficiency of the support, simulated as a translational spring boundary condition with stiffness $k_{\mathrm{v}}$. This quantity can span several orders of magnitude, and therefore we parametrize the model using $\log k_{\mathrm{v}}$ instead. The beam is equipped with $d_{\mathrm{x}}=32$ sensors measuring the vertical displacement, equally spaced along the length as shown in Figure 1(a).